OpenStudy (anonymous):

Pizza delivery times at Pizza Time are normally distributed with a mean time of 27 minutes and a standard deviation of 3 minutes. Approximately what percent of pizzas are delivered between 24 and 30 minutes? 32% 68% 95% 99.7%

OpenStudy (anonymous):

@AG23 @nincompoop @horsegirl27 Can you please help me understand how to do these types of problems.?

OpenStudy (cbarredo1):

Have you studied the Empirical Rule?

OpenStudy (anonymous):

I have not.

OpenStudy (cbarredo1):

A statistical rule stating that for a normal distribution, almost all data will fall within three standard deviations of the mean. Broken down, the empirical rule shows that 68% will fall within the first standard deviation, 95% within the first two standard deviations, and 99.7% will fall within the first three standard deviations of the mean.

OpenStudy (cbarredo1):

ok so we have to figure out how many standard deviations that given interval is from the mean the given interval is [24,30] so how many standard deviations are those 2 values from the mean, which is 27?

OpenStudy (anonymous):

I'm sorry, but what is a standard deviation.? I Quit school for a few purposes and here i am 3 years later and i just need a refreshing. Isn't that kind of like the mean of a data set.?

OpenStudy (cbarredo1):

well 1 std dev = 3 so those values are 3 units away from the mean, right? and 1 std dev = 3 units so we want k here --- [24, 30] = [27-3k, 27+3k] so we can solve either of these equations 24 = 27 - 3k or 30 = 27 + 3k so what is K? So that interval is considered "within 1 standard deviation of the mean"..ok? so that's what we use to get our answer from the empirical rule - A statistical rule stating that for a normal distribution, almost all data will fall within three standard deviations of the mean. Broken down, the empirical rule shows that 68% will fall within the first standard deviation, 95% within the first two standard deviations, and 99.7% will fall within the first three standard deviations of the mean. within the first std dev is the same as within 1 std dev 68% will fall within the first standard deviation that would be [27-3(3), 27+3(3)] within 3 std deviations this is within 1 standard deviation of the mean if the given interval was [18, 36] then it would be 99.7%

OpenStudy (cbarredo1):

The answer would be 68%

OpenStudy (cbarredo1):

It is a quantity calculated to indicate the extent of deviation for a group as a whole.

OpenStudy (anonymous):

okay sweet thank you (:

OpenStudy (cbarredo1):

Did you get it?

OpenStudy (anonymous):

I understand it now.